# Magnetic field measurements in a sample of Class I and Flat-Spectrum protostars observed with SPIRou

### 18 Mar 2026 v1

@@ 3 Analysis and results
In this section, we present the results of our analysis performed on the photospheric lines of the spectra, both in intensity and polarisation, with the final objective being to detect the Zeeman signature in the Stokes $V$ spectra, and to determine either the magnetic strength, or an upper limit on the dipolar magnetic field when no magnetic field is detected. We detected photospheric lines in the spectra of all our sources, except for one: the FS protostar IRAS 03413+3202. Furthermore, the signal-to-noise ratio (S/N) values obtained for two Class I sources with H magnitude greater than 10 (GY92 214 and GY92 378) are too low to constrain either their effective temperatures or their magnetic fields. In the following, we will therefore exclude those three objects. In addition, the dataset of V347 Aur has already been analysed and published in Donati et al. (2024a). The methods used in this paper are similar to those we have used on our data. We have therefore not re-done the analysis for this object, but we kept the source in our sample for the statistical analysis we performed and in the discussion. We present below the photospheric line analysis we performed on the 14 remaining sources.
@@ 3.1 Effective temperature determination
Each spectrum was visually inspected, and only the fits obtained from the highest quality data were retained. When this method did not yield a reliable solution, due to excessive line broadening, low S/N, or strong veiling, we adopted another method. We have computed synthetic spectra using the MARCS atmospheric models and SYNTH3 radiative transfer code. We used the Vienna Atomic Line Database [^4] https://vald.astro.uu.se/ (VALD) to compute spectra over the SPIRou wavelength range. We have adopted microturbulent and macroturbulent velocities of 2 km s-1, solar abundances, and a surface gravity of $\log g=3.5$ (cgs). We have computed a grid of models of effective temperatures between 3500 and 6000 K with a step of 250 K, and of $v\sin i$ values from 5 to 80 km s-1 with a step of 5 km s-1. For each observation we have selected one or two spectral regions in the H band, which gather spectral lines of enough S/N, and sensitive to the effective temperature. We used the independently determined veiling and radial velocity values, and then computed a $\chi^{2}$ value for each model. The effective temperature was selected based on the $\chi^{2}$ minimisation criterion. This method has the advantage to not fit too many parameters that cannot be constrained with the quality of our data, but to constrain the temperature with uncertainties around 250 K, which is enough for the computation of the Least Squares Deconvolution (LSD) profiles (see Section 3.2). Indeed, a variation of $\pm$500 K of the mask does not significantly change the LSD profile shapes, the $B_{\ell}$ values, and the corresponding uncertainties. We emphasise that the quality of our data does not allow us to constrain the microturbulent velocity, and we therefore choose to fix it to usual values for young cool stars, around 1 or 2 km s-1 (Doppmann et al., 2005). As the broadening of the spectral lines is dominated by $v\sin i$ (see Table 5), the choice of the microturbulent velocity value has a negligible effect on the temperature fit. Similarly, the macroturbulent velocity cannot be constrained with our data. Indeed, it affects only the wings of the spectral lines, which are usually lost in the noise. Our choices of 0 to 2 km s-1 for the macroturbulent velocity have a negligible impact on the effective temperature determination.
@@
Determining the effective temperature of VV CrA SW is challenging due to the weakness of its photospheric lines. According to Prato et al. (2003), the SW component does not differ significantly from the NE component. However, the color-color diagram (see Fig. 3 in Prato et al. 2003) suggests a slightly lower temperature. VV CrA SW is therefore expected to be in the range of 3500 and 4000 K. We adopt an intermediate value of 3750 K for the effective temperature. This choice has only a minor impact on the LSD profiles and remains consistent with the presence of photospheric features observed in the LSD profiles (see Appendix E), indicating that the corresponding temperature is appropriate.
@@
Table 1: Effective temperatures of our sample  [^6] Effective temperatures ($T_{\rm eff}$) and their uncertainties ($\sigma$) determined using either ZeeTurbo or SYNTH3.
@@ 3.2 Least-Squares Deconvolved profiles
The LSD is a multi-line technique for computing a pseudo-average line profile as a function of velocity, from a spectrum (Donati et al., 1997). By combining the signal contained in all the spectral lines simultaneously, this method allows us to significantly increase the S/N of our data. The circular polarisation induced by the Zeeman effect at the surface of protostars is expected to be of the order of a few $10^{-4}$ times the continuum intensity, for a typical average line-of-sight magnetic strength of a few hundred Gauss, in a moderate rotator ($v\sin i\sim$20 km s-1). This signal cannot be detected in individual lines of spectra, even with S/N of 100 to 200. The LSD method allows gaining a factor of a few tens in the S/N (depending on the spectral type, and on the level of circumstellar contamination in the spectra), enabling the detection of stellar magnetic fields.
@@ 3.3 Magnetic field detection
We can observe in Figure 8, that the $N$ profiles are consistent with zero, confirming that the polarimeter and data reduction perform as expected. If the $V$ profile shows a significant non-zero signal, while the $N$ profile remains null, we are confident that this signature is of stellar origin.
@@ 3.4 Mean longitudinal magnetic field ($B_{\ell}$)
$B_{\ell}=-2.14\times 10^{11}\frac{\int\frac{V(v)}{I_{c}}vdv}{\lambda zc\int\left[1-\frac{I(v)}{I_{c}}\right]dv}$ (1)
where $\lambda$ is the central wavelength of the spectral line in nm, $z$ is the Landé factor, $c$ is the speed of light in vacuum, $v$ is the velocity, $V$ and II are the Stokes parameters, and $I_{\rm c}$ is the intensity of the continuum. We apply this equation to LSD profiles. We have, therefore, used the normalisation values from the LSD computations: 1660 nm for $\lambda$ and 1.2 for $z$. We used the same velocity range as used for the FAP determination (Table 3). The LSD profiles are not shifted to the stellar rest frame, resulting in asymmetric velocity windows. The uncertainties on $B_{\ell}$ are computed by propagating the errors of the $I$ and $V$ profiles in Eq. 1. The values of $B_{\ell}$ are listed in Table 5. They range from $\sim$80 G to $\sim$200 G in absolute value for the five magnetic protostars.
@@ 4.1 Study of observational bias
For some parameters such as $T_{\rm eff}$, $v\sin i$, and H mag, the tests are performed per source, while for the other parameters (e.g. veiling, $\sigma$($B_{\ell}$), S/N) they are performed per observation. The source sample contains the 14 protostars analysed in Sect. 3, to which we added V347 Aur, analysed in Donati et al. (2024a), for which we used the observation of November 22nd 2022, representative of the 80 observations obtained for that source. The observation sample contains all observations listed in Table 4 plus the V347 Aur observation. The veiling is not available for VV CrA SW because of the weak photospheric lines. The results of the tests are summarized in Table 2.
@@
Table 2: Results of the KS tests between the ND and DD samples [^8] From the top to bottom the tested parameters are the effective temperature, projected rotational velocity, veiling in the H- and K- bands, uncertainty on $B_{\ell}$, the S/N in the spectrum and in the LSD profiles, and the magnitude in H. The $s$ and $p$ values of the KS tests are given between ND and $D$. Bold fonts highlight the tests that significantly reject the null hypothesis at a 95% confidence level (i.e., $p<0.05$).
@@
Our sample exhibits relatively simple Zeeman signatures, suggesting that the magnetic geometry is dominated by low order large-scale magnetic field. The rotational broadening of the spectral lines could also affect magnetic detectability. Indeed, for a similar magnetic field, the Stokes $V$ signature in a star with a low $v\sin i$ will have a larger amplitude than in a star with a large $v\sin i$, as the same Zeeman signature will be spread over a smaller velocity range in the first case compared to the second one (presuming that the large-scale field is simple enough and regions of opposite polarities do not cancel out their respective signatures). If the S/N is the same in both cases, then the signal in the second case will be more diluted in the noise, and the FAP will be higher. When $v\sin i$ is larger, the magnetic detection probability is therefore lower.
Whatever the sample considered, we find that the null hypothesis cannot be rejected. Therefore, no significant difference is found between the DD and ND $v\sin i$ distributions. Remarkably, we are even able to detect magnetic fields in stars with $v\sin i$ as high as $\sim$55 km s-1. At the same time, large-scale magnetic fields are not detected in many protostars with low $v\sin i$ (Fig. 3). Because, for almost all stars in our sample, the $v\sin i$ was unknown or poorly constrained at the time of the observations, we could not take this parameter into account in the computation of our exposure times. Our observational strategy appears, therefore, adequate as it allowed to avoid a potential $v\sin i$ observational bias.
Another parameter that may influence magnetic detections is the veiling. Veiling refers to the dilution of photospheric absorption lines due to excess continuum emission, typically interpreted as a signature of active accretion in young stars (Connelley and Greene, 2010). If a star exhibits strong veiling, this can affect the detectability of magnetic fields by reducing the line contrast with respect to the intensity continuum and, therefore, increasing the relative noise in Stokes $I$ and $V$ profiles (see Appendix C). The KS tests reject the null hypothesis at a confidence level higher than 95% when comparing the veiling distributions of the DD and ND samples, in both H and K bands, and in both statistical samples. In Fig. 4, we indeed find that the higher the veiling, the lower the detection rate.
@@
Finally, we may also wonder whether the quality of our data affects magnetic detection. We find no significant differences between the S/N distributions of the DD and ND samples, considering the S/N in either the spectra (S/Ns), or the LSD profiles (S/NLSD).
Our magnetic detection rate is not affected by the S/N of our data but is affected by veiling. This means that we should observe a correlation between the magnetic field detections and the uncertainties of the longitudinal field measurements (see Appendix C). Indeed, we find that the distributions of $\sigma$($B_{\ell}$) for the DD and ND samples are significantly different (Table 2). Furthermore, Fig. 4 shows that the uncertainties on $B_{\ell}$ are systematically lower for the least veiled spectra. The veiling phenomenon could therefore be the main source of difficulty for detecting magnetic fields in Class I and FS sources.
@@ 4.2 Upper limits on the dipole magnetic field
Stars in which we did not detect a magnetic field may still host one at their surface, but it may be too weak or too complex to be detected with our current data quality. In this section, we aim to estimate the maximal magnetic field strength a star can have that would still remain hidden in the noise of our data. In the case of a complex magnetic field, this is not easy because such fields can exhibit a large variety of topologies, and a large amplitude of strength, which could remain hidden within a given noise level. On the other hand, magnetic fields observed in fully convective stars (i.e. in CTTS or M-dwarfs) are mainly dipolar (Donati et al., 2011; Gregory et al., 2012; Morin et al., 2008, 2010). Furthermore, our sample mostly contains simple Zeeman signatures, most likely produced by low-order magnetic fields. We will therefore only consider a dipole to compute the upper limit.
@@
Table 3: Velocity range and upper limits on the dipolar magnetic field derived using the marginal (MD) and definite ($D$) detection criteria for each star [^9] The detected magnetic protostars are indicated in bold font.
@@ VV CrA NE 25/20-25/20 4700 ¿5000
To assess the reliability of these upper limit values, we also computed them for stars with detected magnetic fields. We find values between approximately 2 kG and more than 5 kG. We observe that the highest upper limits can be found independently in detected and non-detected magnetic protostars. The projected rotational velocity affects the upper limit on the dipole field strength. A higher $v\sin i$ dilutes the Stokes $V$ signature, implying that stronger magnetic fields are required for a detection in stars with large $v\sin i$, compared to stars with low $v\sin i$. We find, indeed, that all sources with a limit above 5 kG have a $v\sin i$ larger than 35 km s-1. The fact that such large limits are found in DD and ND protostars means that large $v\sin i$ is not the main parameter that hampers the detection of magnetic field. This is in line with the similar $v\sin i$ distributions that we find between the DD and ND samples (Sect. 4.1).
@@ 4.3 Impact of our magnetic measurements on the origin and evolution of stellar magnetic fields
We conducted a spectropolarimetric analysis of a sample of Class I and Flat-Spectrum protostars, aiming to investigate the presence of large-scale magnetic fields. All young, low-mass observed protostars are expected to be fully convective or to have a large convective envelope (Palla, 1998; Baraffe et al., 2017), we expect all of them to host detectable magnetic fields, generated by convective dynamos, as in low-mass main-sequence stars. However, magnetic fields were detected in only 6 out of the 15 stars of our final sample, corresponding to 40 % of the sample. This relatively low detection rate raises several questions regarding the magnetic origin of these young stars, their internal processes, and their magnetospheric environments. In particular, the absence or apparent non-detection of large-scale magnetic fields in a substantial fraction of Class I and FS objects challenges the assumption that such fields are ubiquitous at these evolutionary stages.
@@
A central question that emerges is whether large-scale magnetic fields, commonly observed in more evolved systems such as CTTS, are truly universal across all largely-convective stellar objects. Assuming that our protostars lie on the birthline, we expect them to be fully convective up to an effective temperature of 4750 K, and have a convective envelope of mass larger than 40 % of the stellar mass up to 5250 K (Villebrun et al., 2019). Under this assumption, $\sim$80 % of our sample is fully convective, while the remainder has a radiative core plus a large convective envelope. We can also estimate the range of mass probed by our sample using the effective temperatures derived in Sect. 3.1. We find a range of 0.2 to 2.5 M. This means that their direct descendants are the CTTS of low- and intermediate-mass (Herbig, 1962; Villebrun et al., 2019). The magnetic properties of these objects have been intensely characterised thanks to dense spectropolarimetric monitoring, and Zeeman Doppler Imaging analysis (e.g. Hussain et al., 2009; Donati et al., 2013; Pouilly et al., 2020; Nowacki et al., 2023; Zaire et al., 2024). Magnetic fields are detected in fully convective stars, and almost all of them host a strong (>> 1 kG) magnetic field, dominated by a dipolar component, and sometimes a multipolar magnetic field. When published, the maximum absolute $B_{\ell}$ values are larger than 400 G and can reach 900 G. Our snapshot observations do not allow us to perform ZDI to retrieve the topology of the magnetic field we have detected in our sample. However, we do not detect strong magnetic fields in all of them: the maximum absolute $B_{\ell}$ value is 200 G. A single measurement of $B_{\ell}$ is strongly dependent on the rotation phase. Our phase sampling is less complete than that of the ZDI studies, therefore we may have missed the rotation phases when $B_{\ell}$ is maximum. However, out of the 16 observations we obtained for the 5 magnetic protostars, none are higher than 200  G, which are at most similar, but mainly weaker than in fully-convective CTTS (Donati et al., 2024b, c; Nowacki et al., 2023).
@@
On one hand, Class I and FS protostars have a larger radius than CTTS as they are less evolved. Class I and FS protostars are younger than 1 Myr, but probably older than 0.5 Myr (Fiorellino et al., 2023). CTTS are 1 to 3 Myr old (e.g. Zaire et al., 2024). According to stellar evolutionary models (e.g. Baraffe et al., 2015), from 0.5 Myr to 3 Myr, the radius decreases by a factor up to 2. If we assume a constant magnetic flux as the protostars evolve towards the pre-main-sequence, then the surface magnetic field can increase by a factor up to 4, which could explain why the $B_{\ell}$ in our sample are in general weaker than at later stages. On the other hand, it is not clear whether magnetic flux conservation can be invoked when magnetic fields are of dynamo origin. Indeed, the major structural difference between Class I and fully-convective Class II stars is the surface gravity, hence the gravity gradient within the convection zone. Magneto-hydro-dynamical models of convective stars of various stratifications are required to understand the role of stratification in the efficiency of dynamo processes. Within the PROMETHEE project, we are developing models for different stratification levels, which will help us to interpret this result (Guseva et al., in prep.).
@@
Finally, it is interesting to mention that in fully convective main-sequence M dwarfs, a phenomenon known as dynamo bistability has been observed (Morin et al., 2010). In this scenario, stars with identical global properties (mass, rotation rate, convection regime) can exhibit either strong, predominantly dipolar magnetic fields or much weaker, complex, and multipolar fields. It is plausible that a similar mechanism is at play in our sample, and could explain part of the non-detections.

### 22 Mar 2026 v2

@@ 3 Analysis and results
In this section we present the results of our analysis performed on the photospheric lines of the spectra, both in intensity and polarisation, with the final objective being to detect the Zeeman signature in the Stokes $V$ spectra, and to determine either the magnetic strength or an upper limit on the dipolar magnetic field when no magnetic field is detected. We detected photospheric lines in the spectra of all our sources, except for one: the FS protostar IRAS 03413+3202. Furthermore, the signal-to-noise ratio (S/N) values obtained for two Class I sources with H magnitude greater than 10 (GY92 214 and GY92 378) are too low to constrain either their effective temperatures or their magnetic fields. In the following, we therefore exclude those three objects. In addition, the dataset of V347 Aur has already been analysed and published in Donati et al. (2024a). The methods used in this paper are similar to those we have used on our data. We have therefore not re-done the analysis for this object, but we kept the source in our sample for the statistical analysis we performed and in the discussion. We present below the photospheric line analysis we performed on the 14 remaining sources.
@@ 3.1 Effective temperature determination
Each spectrum was visually inspected, and only the fits obtained from the highest quality data were retained. When this method did not yield a reliable solution, due to excessive line broadening, low S/N, or strong veiling, we adopted another method. We have computed synthetic spectra using the MARCS atmospheric models and SYNTH3 radiative transfer code. We used the Vienna Atomic Line Database [^4] https://vald.astro.uu.se/ (VALD) to compute spectra over the SPIRou wavelength range. We have adopted microturbulent and macroturbulent velocities of 2 km s-1, solar abundances, and a surface gravity of $\log g=3.5$ (cgs). We have computed a grid of models of effective temperatures between 3500 and 6000 K with a step of 250 K, and of $v\sin i$ values from 5 to 80 km s-1 with a step of 5 km s-1. For each observation we have selected one or two spectral regions in the H band, which gather spectral lines of enough S/N, and sensitive to the effective temperature. We used the independently determined veiling and radial velocity values, and then computed a $\chi^{2}$ value for each model. The effective temperature was selected based on the $\chi^{2}$ minimisation criterion. This method has the advantage to not fit too many parameters that cannot be constrained with the quality of our data, but to constrain the temperature with uncertainties around 250 K, which is enough for the computation of the least-squares deconvolution (LSD) profiles (see Sect. 3.2). A variation of $\pm$500 K of the mask does not significantly change the LSD profile shapes, the $B_{\ell}$ values, and the corresponding uncertainties. We emphasise that the quality of our data does not allow us to constrain the microturbulent velocity, and we therefore choose to fix it to usual values for young cool stars, around 1 or 2 km s-1 (Doppmann et al., 2005). As the broadening of the spectral lines is dominated by $v\sin i$ (see Table 5), the choice of the microturbulent velocity value has a negligible effect on the temperature fit. Similarly, the macroturbulent velocity cannot be constrained with our data. It affects only the wings of the spectral lines, which are usually lost in the noise. Our choices of 0 to 2 km s-1 for the macroturbulent velocity have a negligible impact on the effective temperature determination.
@@
Determining the effective temperature of VV CrA SW is challenging due to the weakness of its photospheric lines. According to Prato et al. (2003), the SW component does not differ significantly from the NE component. However, the colour-colour diagram (see Fig. 3 in Prato et al. 2003) suggests a slightly lower temperature. VV CrA SW is therefore expected to be in the range of 3500 and 4000 K. We adopt an intermediate value of 3750 K for the effective temperature. This choice has only a minor impact on the LSD profiles and remains consistent with the presence of photospheric features observed in the LSD profiles (see Appendix E), indicating that the corresponding temperature is appropriate.
@@
Table 1: Effective temperatures of our sample.  [^6] Effective temperatures ($T_{\rm eff}$) and their uncertainties ($\sigma$) determined using either ZeeTurbo or SYNTH3.
@@ 3.2 Least-squares deconvolved profiles
The LSD is a multi-line technique for computing a pseudo-average line profile as a function of velocity, from a spectrum (Donati et al., 1997). By combining the signal contained in all the spectral lines simultaneously, this method allows us to significantly increase the S/N of our data. The circular polarisation induced by the Zeeman effect at the surface of protostars is expected to be of the order of a few $10^{-4}$ times the continuum intensity, for a typical average line-of-sight magnetic strength of a few hundred gauss, in a moderate rotator ($v\sin i\sim$20 km s-1). This signal cannot be detected in individual lines of spectra, even with S/N of 100 to 200. The LSD method allows an increase of a factor of a few tens in the S/N (depending on the spectral type, and on the level of circumstellar contamination in the spectra), enabling the detection of stellar magnetic fields.
@@ 3.3 Magnetic field detection
We can observe in Fig. 8 that the $N$ profiles are consistent with zero, confirming that the polarimeter and data reduction perform as expected. If the $V$ profile shows a significant non-zero signal, while the $N$ profile remains null, we are confident that this signature is of stellar origin.
@@ 3.4 Mean longitudinal magnetic field ($B_{\ell}$)
$B_{\ell}=-2.14\times 10^{11}\frac{\int\frac{V(v)}{I_{c}}vdv}{\lambda zc\int\left[1-\frac{I(v)}{I_{c}}\right]dv}.$ (1)
Here $\lambda$ is the central wavelength of the spectral line in nm, $z$ is the Landé factor, $c$ is the speed of light in vacuum, $v$ is the velocity, $V$ and II are the Stokes parameters, and $I_{\rm c}$ is the intensity of the continuum. We apply this equation to LSD profiles. We have therefore used the normalisation values from the LSD computations: 1660 nm for $\lambda$ and 1.2 for $z$. We used the same velocity range as used for the FAP determination (Table 3). The LSD profiles are not shifted to the stellar rest frame, resulting in asymmetric velocity windows. The uncertainties on $B_{\ell}$ are computed by propagating the errors of the $I$ and $V$ profiles in Eq. 1. The values of $B_{\ell}$ are listed in Table 5. They range from $\sim$80 G to $\sim$200 G in absolute value for the five magnetic protostars.
@@ 4.1 Study of observational bias
For some parameters such as $T_{\rm eff}$, $v\sin i$, and H mag, the tests are performed per source, while for the other parameters (e.g. veiling, $\sigma$($B_{\ell}$), S/N) they are performed per observation. The source sample contains the 14 protostars analysed in Sect. 3, to which we added V347 Aur, analysed in Donati et al. (2024a), for which we used the observation of November 22nd 2022, representative of the 80 observations obtained for that source. The observation sample contains all observations listed in Table 4 plus the V347 Aur observation. The veiling is not available for VV CrA SW because of the weak photospheric lines. The results of the tests are summarised in Table 2.
@@
Table 2: Results of the KS tests between the ND and DD samples. [^8] From top to bottom the tested parameters are the effective temperature, projected rotational velocity, veiling in the H- and K- bands, uncertainty on $B_{\ell}$, the S/N in the spectrum and in the LSD profiles, and the magnitude in H. The $s$ and $p$ values of the KS tests are given between ND and $D$. Bold fonts highlight the tests that significantly reject the null hypothesis at a 95% confidence level (i.e. $p<0.05$).
@@
Our sample exhibits relatively simple Zeeman signatures, suggesting that the magnetic geometry is dominated by low-order large-scale magnetic field. The rotational broadening of the spectral lines could also affect magnetic detectability. For a similar magnetic field, the Stokes $V$ signature in a star with a low $v\sin i$ will have a larger amplitude than in a star with a large $v\sin i$, as the same Zeeman signature will be spread over a smaller velocity range in the first case compared to the second one (presuming that the large-scale field is simple enough and regions of opposite polarities do not cancel out their respective signatures). If the S/N is the same in both cases, then the signal in the second case will be more diluted in the noise, and the FAP will be higher. When $v\sin i$ is larger, the magnetic detection probability is therefore lower.
Whatever the sample considered, we find that the null hypothesis cannot be rejected. Therefore, no significant difference is found between the DD and ND $v\sin i$ distributions. Remarkably, we are even able to detect magnetic fields in stars with $v\sin i$ as high as $\sim$55 km s-1. At the same time, large-scale magnetic fields are not detected in many protostars with low $v\sin i$ (Fig. 3). Because, for almost all stars in our sample, the $v\sin i$ was unknown or poorly constrained at the time of the observations, we could not take this parameter into account in the computation of our exposure times. Our observational strategy therefore appears adequate as it allowed a potential $v\sin i$ observational bias to be avoided.
Another parameter that may influence magnetic detections is the veiling. Veiling refers to the dilution of photospheric absorption lines due to excess continuum emission, typically interpreted as a signature of active accretion in young stars (Connelley and Greene, 2010). If a star exhibits strong veiling, this can affect the detectability of magnetic fields by reducing the line contrast with respect to the intensity continuum, and therefore increasing the relative noise in Stokes $I$ and $V$ profiles (see Appendix C). The KS tests reject the null hypothesis at a confidence level higher than 95% when comparing the veiling distributions of the DD and ND samples, in both H and K bands, and in both statistical samples. In Fig. 4, we indeed find that the higher the veiling, the lower the detection rate.
@@
Finally, we may also wonder whether the quality of our data affects magnetic detection. We find no significant differences between the S/N distributions of the DD and ND samples, considering the S/N in either the spectra (S/Ns) or the LSD profiles (S/NLSD).
Our magnetic detection rate is not affected by the S/N of our data but is affected by veiling. This means that we should observe a correlation between the magnetic field detections and the uncertainties of the longitudinal field measurements (see Appendix C). We find that the distributions of $\sigma$($B_{\ell}$) for the DD and ND samples are significantly different (Table 2). Furthermore, Fig. 4 shows that the uncertainties on $B_{\ell}$ are systematically lower for the least veiled spectra. The veiling phenomenon could therefore be the main source of difficulty for detecting magnetic fields in Class I and FS sources.
@@ 4.2 Upper limits on the dipole magnetic field
Stars in which we did not detect a magnetic field may still host one at their surface, but it may be too weak or too complex to be detected with our current data quality. In this section we aim to estimate the maximal magnetic field strength a star can have that would still remain hidden in the noise of our data. In the case of a complex magnetic field, this is not easy because such fields can exhibit a large variety of topologies, and a large amplitude of strength, which could remain hidden within a given noise level. On the other hand, magnetic fields observed in fully convective stars (i.e. in CTTSs or M dwarfs) are mainly dipolar (Donati et al., 2011; Gregory et al., 2012; Morin et al., 2008, 2010). Furthermore, our sample mostly contains simple Zeeman signatures, most likely produced by low-order magnetic fields. We therefore only consider a dipole to compute the upper limit.
@@
Table 3: Velocity range and upper limits on the dipolar magnetic field derived using the marginal (MD) and definite ($D$) detection criteria for each star. [^9] The detected magnetic protostars are indicated in bold font.
@@ VV CrA NE 25/20-25/20 4700 ¿5000
To assess the reliability of these upper limit values, we also computed them for stars with detected magnetic fields. We find values between approximately 2 kG and more than 5 kG. We observe that the highest upper limits can be found independently in detected and non-detected magnetic protostars. The projected rotational velocity affects the upper limit on the dipole field strength. A higher $v\sin i$ dilutes the Stokes $V$ signature, implying that stronger magnetic fields are required for a detection in stars with large $v\sin i$, compared to stars with low $v\sin i$. We find that all sources with a limit above 5 kilogauss have a $v\sin i$ larger than 35 km s-1. The fact that such large limits are found in DD and ND protostars means that large $v\sin i$ is not the main parameter that hampers the detection of magnetic field. This is in line with the similar $v\sin i$ distributions that we find between the DD and ND samples (Sect. 4.1).
@@ 4.3 Impact of our magnetic measurements on the origin and evolution of stellar magnetic fields
We conducted a spectropolarimetric analysis of a sample of Class I and FS protostars, aiming to investigate the presence of large-scale magnetic fields. All young, low-mass observed protostars are expected to be fully convective or to have a large convective envelope (Palla, 1998; Baraffe et al., 2017), we expect all of them to host detectable magnetic fields, generated by convective dynamos, as in low-mass main-sequence stars. However, magnetic fields were detected in only 6 out of the 15 stars of our final sample, corresponding to 40 % of the sample. This relatively low detection rate raises several questions regarding the magnetic origin of these young stars, their internal processes, and their magnetospheric environments. In particular, the absence or apparent non-detection of large-scale magnetic fields in a substantial fraction of Class I and FS objects challenges the assumption that such fields are ubiquitous at these evolutionary stages.
@@
A central question that emerges is whether large-scale magnetic fields, commonly observed in more evolved systems such as CTTSs, are truly universal across all largely convective stellar objects. Assuming that our protostars lie on the birthline, we expect them to be fully convective up to an effective temperature of 4750 K, and have a convective envelope of mass larger than 40 % of the stellar mass up to 5250 K (Villebrun et al., 2019). Under this assumption, $\sim$80 % of our sample is fully convective, while the remainder has a radiative core plus a large convective envelope. We can also estimate the range of mass probed by our sample using the effective temperatures derived in Sect. 3.1. We find a range of 0.2 to 2.5 M. This means that their direct descendants are the CTTSs of low- and intermediate-mass (Herbig, 1962; Villebrun et al., 2019). The magnetic properties of these objects have been intensely characterised thanks to dense spectropolarimetric monitoring, and ZDI analysis (e.g. Hussain et al., 2009; Donati et al., 2013; Pouilly et al., 2020; Nowacki et al., 2023; Zaire et al., 2024). Magnetic fields are detected in fully convective stars, and almost all of them host a strong (>> 1 kG) magnetic field, dominated by a dipolar component, and sometimes a multipolar magnetic field. When published, the maximum absolute $B_{\ell}$ values are larger than 400 G and can reach 900 G. Our snapshot observations do not allow us to perform ZDI to retrieve the topology of the magnetic field we have detected in our sample. However, we do not detect strong magnetic fields in all of them: the maximum absolute $B_{\ell}$ value is 200 G. A single measurement of $B_{\ell}$ is strongly dependent on the rotation phase. Our phase sampling is less complete than that of the ZDI studies, and therefore we may have missed the rotation phases when $B_{\ell}$ is maximum. However, out of the 16 observations we obtained for the 5 magnetic protostars, none are higher than 200  G, which are at most similar, but mainly weaker than in fully convective CTTSs (Donati et al., 2024b, c; Nowacki et al., 2023).
@@
On one hand, Class I and FS protostars have a larger radius than CTTSs as they are less evolved. Class I and FS protostars are younger than 1 Myr, but probably older than 0.5 Myr (Fiorellino et al., 2023). CTTSs are 1 to 3 Myr old (e.g. Zaire et al., 2024). According to stellar evolutionary models (e.g. Baraffe et al., 2015), from 0.5 Myr to 3 Myr, the radius decreases by a factor up to 2. If we assume a constant magnetic flux as the protostars evolve towards the pre-main sequence, then the surface magnetic field can increase by a factor up to 4, which could explain why the $B_{\ell}$ in our sample are in general weaker than at later stages. On the other hand, it is not clear whether magnetic flux conservation can be invoked when magnetic fields are of dynamo origin. The major structural difference between Class I and fully convective Class II stars is the surface gravity, hence the gravity gradient within the convection zone. Magneto-hydro-dynamical models of convective stars of various stratifications are required to understand the role of stratification in the efficiency of dynamo processes. Within the PROMETHEE project, we are developing models for different stratification levels that will help us to interpret this result (Guseva et al., in prep.).
@@
Finally, it is interesting to mention that in fully convective main-sequence M dwarfs, a phenomenon known as dynamo bistability has been observed (Morin et al., 2010). In this scenario, stars with identical global properties (mass, rotation rate, convection regime) can exhibit either strong predominantly dipolar magnetic fields or much weaker, complex, and multipolar fields. It is plausible that a similar mechanism is at play in our sample, and could explain part of the non-detections.